OpenStudy (anonymous):

The numerator of a fraction is 3 times the denominator. If 5 is subtracted from the numerator, and added to the denominator, the resulting fraction is equal to 16/12

OpenStudy (anonymous):

16+5=21 12-5=7 21/7 is the answer. Check: 3*7=21